# Terrestrial Matter Effects on Reactor Antineutrino Oscillations: Constant vs. Fluctuated Density Profiles

### v1

@@ Acknowledgments
Jing-yu Zhu thanks Wen-jie Wu for helpful discussions. The work of JYZ was supported in part by the National Key Research and Development Program of China (under grant number 2021YFA1601300) and by the National Natural Science Foundation of China (under grant number 12505134). The work of YFL was supported in part by the National Natural Science Foundation of China (under grant numbers 12075255). The work of ADW was supported in part by the Open Project Program of National Key Laboratory of Uranium Resources Exploration-Mining and Nuclear Remote Sensing (under grant number NKLUR-2024-QN-005). The work of YX was supported in part by the National Natural Science Foundation of China (under grant number 42074092).

### v3

@@ Acknowledgments
Jing-yu Zhu thanks Wen-jie Wu for helpful discussions. The work of Jing-yu Zhu was supported in part by the National Key Research and Development Program of China (under grant number 2021YFA1601300) and by the National Natural Science Foundation of China (under grant number 12505134). The work of Yu-Feng Li was supported in part by the National Natural Science Foundation of China (under grant number 12075255). The work of Andong Wang was supported in part by the Open Project Program of National Key Laboratory of Uranium Resources Exploration-Mining and Nuclear Remote Sensing (under grant number NKLUR-2024-QN-005). The work of Ya Xu was supported in part by the National Natural Science Foundation of China (under grant number 42074092).